OpenStudy (anonymous):

7 3/4 - 3/8 = 7 1/4 7 1/8 7 3/8 7 1/2 Not Here

OpenStudy (anonymous):

7.75−0.375=7.375=7 3/8

OpenStudy (anonymous):

it is (73/4) -(3/8) =(143/8)